# Challenges and proposed solutions in modeling multimodal data: A systematic review

### v3

@@ 2.1.5 Optimal fusion technique
In addressing the challenge of finding the optimal fusion strategy, the problem in the context of deep multimodal learning can be reduced to two central architectural design choices. The first choice relates to the level at which different modalities should be combined, which can be guided by considering the challenges and advantages of early, late, and intermediate fusion approaches. The second architectural design choice concerns which modalities to combine, as including all available modalities does not always enhance the performance of a machine learning algorithm 101. A total of 13 studies in our review identified finding the optimal fusion strategy as a challenge. For example, Xu et al. 102 tackled this challenge by extending the neural architecture search 103, 104, which was originally designed for unimodal data, to incorporate the joint optimization of fusion strategies and modality-specific architectures for multimodal data. This extension allows for searching for unique architectures for distinct modalities and identifies the best strategies to fuse these architectures at the right representation level, and then they utilize state-of-the-art neural architecture search methods to explore and optimize fusion strategies for multimodal data. By jointly optimizing fusion strategies and modality-specific architectures, their approach aims to enhance prediction performance on EHR tasks while controlling for computation costs.
@@
In the fusion of imaging and clinical data, Wang et al. 30 pointed out a common heuristic approach to data fusion in many studies, where data from multiple sources are simply concatenated without proper preprocessing or assessment of fusion quality. This failure to interpret the fused feature space can result in suboptimal model performance, particularly in medical imaging problems where the features of different data sources can vary significantly. Therefore, they proposed modeling the fully connected layer of a deep neural network as a potential function following the classical Gibbs measure. By considering the features of the fully connected layer as random variables governed by state functions representing different data sources, the method quantifies source bias and aims to optimize data fusion quality. The use of a vector-growing encoding scheme, such as positional encoding, helps the conversion of low-dimensional clinical features into a rich feature space to complement high-dimensional imaging features, ultimately enhancing the fusion process in multi-source deep learning problems.
@@
Several studies have highlighted that different modalities often have unequal importance and contributions, necessitating innovative methods to adaptively weight modalities during the fusion process. For instance, Dai et al. 19 employs dual encoders and joint attention mechanisms to assign greater importance to features that are most relevant for Alzheimer’s disease classification. Similarly, Hatami et al. 38 introduces a weighted ensemble method for predicting stroke outcomes, where weights are assigned to imaging module predictions based on their alignment with clinical metadata, thereby improving prediction accuracy. In another approach, Rahaman et al. 26 uses attention mechanisms to prioritize the most informative modalities, mitigating performance degradation that arises from forcing equal contributions. Lastly, Xia et al. 31 incorporates an adaptive fusion layer to dynamically assign modality-specific weights, optimizing their contributions in clustering tasks. Collectively, these studies underscore the critical role of adaptive weighting techniques in enhancing the effectiveness of multimodal fusion by accounting for the varying importance of different modalities. In the classification task, there are various approaches to determining the optimal fusion technique. For example, Pawłowski et al. 5 emphasized the significance of selecting the most appropriate fusion technique. Specifically in classification tasks, they outlined three criteria for making design decisions when selecting the most suitable data fusion technique. These criteria involve evaluating the influence of each modality on the machine learning problem, considering the task type, and assessing memory usage during both training and prediction phases.
@@
There are also two similar works by Khan et al. 16 and Jabeen et al. 15 which determine the optimal fusion technique in two steps. Firstly, they select the best features using the differential evolution selected features. Khan et al. 16 introduced a probability-based fusion method that involves computing probabilities for selected features and then choosing one feature based on the highest probability value. This selected feature is used as a reference for comparison, and features from different vectors are merged into a single matrix. This comparison step helps address the issue of redundant features in the vectors. On the other hand, Jabeen et al. 15 proposed the maximum correlation-based fusion approach, aiming to combine optimal feature vectors by selecting features with strong correlations, discarding those with weak correlations, and merging them into a single vector. By prioritizing features with high correlation coefficients, this fusion method enhances the representational power of the fused vector. This approach ensures that the combined features effectively capture complementary information, leading to improved classification accuracy.
2.1.6 Other challenges
@@
Beyond the main challenges detailed in this manuscript, there is a range of unique additional issues that do not align with predefined categories yet are still critical for understanding and improving multimodal data integration. In LABEL:Tab:challenges, the specific challenges for “Other Challenges” are explicitly stated alongside their corresponding solutions, as they vary across different studies. One major challenge is extracting high-order interactions within and among data modalities 55, 78, which is crucial for identifying complex, non-linear relationships that may exist between features. These interactions are particularly relevant in clinical contexts where modalities such as genomics and imaging provide complementary insights. Similarly, capturing the intricate relationships and temporal dynamics across modalities 53, 62, while accounting for irregularities in patient visit timings, is a significant hurdle. Patient data often exhibits temporal patterns, with varying time gaps between visits, making it essential to develop models capable of handling such inconsistencies. High dimensionality adds another layer of complexity 54, 65 as many modalities, such as imaging and genomics, have a vast number of features. This can lead to intensive computational requirements and the risk of overfitting, necessitating robust dimensionality reduction and feature selection techniques. Capturing intermodality interactions 19, 21, 37, 53, 58, 70, 73 further complicates modeling, as it requires mechanisms that can integrate and prioritize information from heterogeneous sources without losing critical cross-modal correlations. Lastly, heterogeneity in data 42, 43, 52, 55, 63, 76, 80, whether in terms of formats, scales, or quality, presents a persistent challenge, particularly in ensuring that diverse modalities are integrated cohesively. These challenges highlight the need for innovative approaches, such as advanced attention mechanisms, hybrid fusion strategies, and efficient computational frameworks, to enable robust and interpretable multimodal data integration in clinical applications.
3 Discussion
@@
A notable aspect of this review is the variety of innovative solutions proposed to address the key challenges in multimodal modeling. For missing data modalities, researchers have introduced methods such as multitask learning 43, matrix completion 79, 87, variational autoencoders 23, 90, generative adversarial networks 51, 71, hybrid graph convolutional networks 20, and knowledge distillation wang_multi-modal_2023. These approaches aim to impute or compensate for absent modalities and have shown improvements in handling incomplete datasets. To overcome limited sample sizes (“small data”), strategies such as data augmentation 25, 15, transfer learning from large pretrained models 29, 22, 15, 41, 59, 50, and knowledge distillation 56 have been employed to maximize the value of scarce data. However, the effectiveness of these techniques in clinical contexts is constrained: many rely on external pretrained models or large datasets that are not always available, and they may fail to capture complex cross-modal patterns when data are sparse or heterogeneous. For improving interpretability, studies have applied attention mechanisms, gradient-based attribution tools 21, 22 , and explanation techniques such as DeepSHAP 42, 44, xu_explainable_2021, 46, which help clarify model decision-making. To address imbalances in the dimensionality of different modalities, where one modality may dominate due to feature richness, methods like feature selection, dimensionality reduction 80, and weighted loss functions 39, 25, 58, 22 have been used to achieve more balanced integration. Finally, the selection of an optimal fusion strategy remains an open challenge. Early, intermediate, late, and hybrid fusion each offer distinct strengths and weaknesses, and recent advances such as neural architecture search and attention-based fusion 19, 26, 45 show promise in tailoring integration methods to specific tasks. Taken together, these findings highlight that there is no one-size-fits-all solution: each challenge requires targeted techniques, and many multimodal pipelines combine multiple methods to achieve robust performance.
@@
Beyond methodological issues, broader concerns of fairness, ethics, and deployment are increasingly recognized as critical in multimodal AI for medicine. Recent reviews emphasize that fairness research is still scarce in many clinical domains, with sensitive attributes such as race, sex, and socio-economic status often underrepresented in available datasets, thereby limiting equity in model performance liu2023translational, norori2023fairness. In medical imaging, for example, studies have shown that algorithms can exploit demographic shortcuts that yield good results on benchmark datasets but fail to generalize to new populations yang2024limits. The multimodal setting introduces additional complexity, as different modalities may carry distinct sources of bias, which can interact or even reinforce each other hickmon2024multimodal. Beyond fairness, deployment in clinical practice raises ethical and practical concerns including data privacy, informed consent for the integration of heterogeneous data sources, interpretability of multimodal decisions for clinicians, and alignment with regulatory frameworks. While addressing these aspects was beyond the scope of our systematic review, they represent important directions for future work to ensure that technical advances in multimodal modeling translate into safe, equitable, and clinically useful applications.
@@ 4 Conclusion
Data Availability Statement
@@
Acknowledgments
Conflict of Interest Statement
The authors declare that the research was conducted in the absence of any commercial or financial relationships that could be construed as a potential conflict of interest.
Generative AI statement
The authors declare that no Gen AI was used in the creation of this manuscript.
Supplemental Material

### v4

@@ Results
In addressing the challenge of finding the optimal fusion strategy, the problem in the context of deep multimodal learning can be reduced to two central architectural design choices. The first choice relates to the level at which different modalities should be combined, which can be guided by considering the challenges and advantages of early, late, and intermediate fusion approaches. The second architectural design choice concerns which modalities to combine, as including all available modalities does not always enhance the performance of a machine learning algorithm Ramachandram and Taylor [2017]. A total of 13 studies in our review identified finding the optimal fusion strategy as a challenge. For example, Xu et al. [2021] tackled this challenge by extending the neural architecture search Yao et al. [1999], Real et al. [2019], which was originally designed for unimodal data, to incorporate the joint optimization of fusion strategies and modality-specific architectures for multimodal data. This extension allows for searching for unique architectures for distinct modalities and identifies the best strategies to fuse these architectures at the right representation level, and then they utilize state-of-the-art neural architecture search methods to explore and optimize fusion strategies for multimodal data. By jointly optimizing fusion strategies and modality-specific architectures, their approach aims to enhance prediction performance on EHR tasks while controlling for computation costs.
@@
In the fusion of imaging and clinical data, Wang et al. [2023] pointed out a common heuristic approach to data fusion in many studies, where data from multiple sources are simply concatenated without proper preprocessing or assessment of fusion quality. This failure to interpret the fused feature space can result in suboptimal model performance, particularly in medical imaging problems where the features of different data sources can vary significantly. Therefore, they proposed modeling the fully connected layer of a deep neural network as a potential function following the classical Gibbs measure. By considering the features of the fully connected layer as random variables governed by state functions representing different data sources, the method quantifies source bias and aims to optimize data fusion quality. The use of a vector-growing encoding scheme, such as positional encoding, helps the conversion of low-dimensional clinical features into a rich feature space to complement high-dimensional imaging features, ultimately enhancing the fusion process in multi-source deep learning problems.
@@
Several studies have highlighted that different modalities often have unequal importance and contributions, necessitating innovative methods to adaptively weight modalities during the fusion process. For instance, Dai et al. [2023] employs dual encoders and joint attention mechanisms to assign greater importance to features that are most relevant for Alzheimer’s disease classification. Similarly, Hatami et al. [2022] introduces a weighted ensemble method for predicting stroke outcomes, where weights are assigned to imaging module predictions based on their alignment with clinical metadata, thereby improving prediction accuracy. In another approach, Rahaman et al. [2023] uses attention mechanisms to prioritize the most informative modalities, mitigating performance degradation that arises from forcing equal contributions. Lastly, Xia et al. [2023] incorporates an adaptive fusion layer to dynamically assign modality-specific weights, optimizing their contributions in clustering tasks. Collectively, these studies underscore the critical role of adaptive weighting techniques in enhancing the effectiveness of multimodal fusion by accounting for the varying importance of different modalities. In the classification task, there are various approaches to determining the optimal fusion technique. For example, Pawłowski et al. [2023] emphasized the significance of selecting the most appropriate fusion technique. Specifically in classification tasks, they outlined three criteria for making design decisions when selecting the most suitable data fusion technique. These criteria involve evaluating the influence of each modality on the machine learning problem, considering the task type, and assessing memory usage during both training and prediction phases.
@@
There are also two similar works by Khan et al. [2020] and Jabeen et al. [2022] which determine the optimal fusion technique in two steps. Firstly, they select the best features using the differential evolution selected features. Khan et al. [2020] introduced a probability-based fusion method that involves computing probabilities for selected features and then choosing one feature based on the highest probability value. This selected feature is used as a reference for comparison, and features from different vectors are merged into a single matrix. This comparison step helps address the issue of redundant features in the vectors. On the other hand, Jabeen et al. [2022] proposed the maximum correlation-based fusion approach, aiming to combine optimal feature vectors by selecting features with strong correlations, discarding those with weak correlations, and merging them into a single vector. By prioritizing features with high correlation coefficients, this fusion method enhances the representational power of the fused vector. This approach ensures that the combined features effectively capture complementary information, leading to improved classification accuracy.
Other challenges
@@
Beyond the main challenges detailed in this manuscript, there is a range of unique additional issues that do not align with predefined categories yet are still critical for understanding and improving multimodal data integration. In LABEL:Tab:challenges, the specific challenges for “Other Challenges” are explicitly stated alongside their corresponding solutions, as they vary across different studies. One major challenge is extracting high-order interactions within and among data modalities Dong et al. [2021], Zhang et al. [2014], which is crucial for identifying complex, non-linear relationships that may exist between features. These interactions are particularly relevant in clinical contexts where modalities such as genomics and imaging provide complementary insights. Similarly, capturing the intricate relationships and temporal dynamics across modalities An et al. [2021], Brand et al. [2020], while accounting for irregularities in patient visit timings, is a significant hurdle. Patient data often exhibits temporal patterns, with varying time gaps between visits, making it essential to develop models capable of handling such inconsistencies. High dimensionality adds another layer of complexity Bahador et al. [2021], Liu et al. [2020] as many modalities, such as imaging and genomics, have a vast number of features. This can lead to intensive computational requirements and the risk of overfitting, necessitating robust dimensionality reduction and feature selection techniques. Capturing intermodality interactions Dai et al. [2023], Krix et al. [2023], Golovanevsky et al. [2022], An et al. [2021], Mao et al. [2021], Alam et al. [2018], Zille et al. [2017] further complicates modeling, as it requires mechanisms that can integrate and prioritize information from heterogeneous sources without losing critical cross-modal correlations. Lastly, heterogeneity in data Rabinovici-Cohen et al. [2022], Saad et al. [2022], Abdelaziz et al. [2021], Dong et al. [2021], Li et al. [2020], Eshaghi et al. [2015], Madabhushi et al. [2011], whether in terms of formats, scales, or quality, presents a persistent challenge, particularly in ensuring that diverse modalities are integrated cohesively. These challenges highlight the need for innovative approaches, such as advanced attention mechanisms, hybrid fusion strategies, and efficient computational frameworks, to enable robust and interpretable multimodal data integration in clinical applications.
Discussion
@@
A notable aspect of this review is the variety of innovative solutions proposed to address the key challenges in multimodal modeling. For missing data modalities, researchers have introduced methods such as multitask learning Saad et al. [2022], matrix completion Thung et al. [2013], Fan et al. [2020], variational autoencoders Martí-Juan et al. [2023], Wu and Goodman [2018], generative adversarial networks Ziegler et al. [2022], Cai et al. [2018], hybrid graph convolutional networks Hou et al. [2023], and knowledge distillation Wang et al. [2023]. These approaches aim to impute or compensate for absent modalities and have shown improvements in handling incomplete datasets. To overcome limited sample sizes (“small data”), strategies such as data augmentation Partin et al. [2023], Jabeen et al. [2022], transfer learning from large pretrained models Sükei et al. [2023], Li et al. [2023], Jabeen et al. [2022], Lim et al. [2022], Shickel et al. [2021], Yoo et al. [2022], and knowledge distillation Guan et al. [2021] have been employed to maximize the value of scarce data. However, the effectiveness of these techniques in clinical contexts is constrained: many rely on external pretrained models or large datasets that are not always available, and they may fail to capture complex cross-modal patterns when data are sparse or heterogeneous. For improving interpretability, studies have applied attention mechanisms, gradient-based attribution tools Krix et al. [2023], Li et al. [2023] , and explanation techniques such as DeepSHAP Rabinovici-Cohen et al. [2022], Sheu et al. [2022], Xu et al. [2021], Wang et al. [2022], which help clarify model decision-making. To address imbalances in the dimensionality of different modalities, where one modality may dominate due to feature richness, methods like feature selection, dimensionality reduction Madabhushi et al. [2011], and weighted loss functions Karaman et al. [2022], Partin et al. [2023], Mao et al. [2021], Li et al. [2023] have been used to achieve more balanced integration. Finally, the selection of an optimal fusion strategy remains an open challenge. Early, intermediate, late, and hybrid fusion each offer distinct strengths and weaknesses, and recent advances such as neural architecture search and attention-based fusion Dai et al. [2023], Rahaman et al. [2023], Vanguri et al. [2022] show promise in tailoring integration methods to specific tasks. Taken together, these findings highlight that there is no one-size-fits-all solution: each challenge requires targeted techniques, and many multimodal pipelines combine multiple methods to achieve robust performance.
@@
Beyond methodological issues, broader concerns of fairness, ethics, and deployment are increasingly recognized as critical in multimodal AI for medicine. Recent reviews emphasize that fairness research is still scarce in many clinical domains, with sensitive attributes such as race, sex, and socio-economic status often underrepresented in available datasets, thereby limiting equity in model performance Liu et al. [2023], Norori et al. [2023]. In medical imaging, for example, studies have shown that algorithms can exploit demographic shortcuts that yield good results on benchmark datasets but fail to generalize to new populations Yang et al. [2024]. The multimodal setting introduces additional complexity, as different modalities may carry distinct sources of bias, which can interact or even reinforce each other Hickmon [2024]. Beyond fairness, deployment in clinical practice raises ethical and practical concerns including data privacy, informed consent for the integration of heterogeneous data sources, interpretability of multimodal decisions for clinicians, and alignment with regulatory frameworks. While addressing these aspects was beyond the scope of our systematic review, they represent important directions for future work to ensure that technical advances in multimodal modeling translate into safe, equitable, and clinically useful applications.
@@ Conclusion
Abbreviations
AI Artificial Intelligence
AUC Area Under the Curve
AUROC Area Under the Receiver Operating Characteristic Curve
AUPRC Area Under the Precision–Recall Curve
CNN Convolutional Neural Network
DL Deep Learning
EHR Electronic Health Record
GAN Generative Adversarial Network
LSTM Long Short-Term Memory
MAE Mean Absolute Error
MAPE Mean Absolute Percentage Error
ML Machine Learning
MRI Magnetic Resonance Imaging
MVAE Multimodal Variational Autoencoder
RMSE Root Mean Square Error
ROC Receiver Operating Characteristic
SVM Support Vector Machine
VAE Variational Autoencoder
Supplementary Information
Acknowledgements
@@
Data Availability
Declarations
Ethics approval and consent to participate
Not applicable
Consent for publication
Not applicable
Competing interests
The authors declare no competing interests